Six men take five days to complete a job.how long would the same job take 15 men.
Answer by stanbon(75887)

Six men take five days to complete a job.
how long would the same job take 15 men.
------
# of men and # of days are inversely related:
d = k/m
------
Solve for "k" using "6 men take 5 days to complete a job":
5 = k/6
k = 30
------
Equation:
d = 30/m
-----
how long would the same job take 15 men.
d = 30/15
days = 2 (# of eays for 15 men to do the job)
-----------
Cheers,
Stan H.
===========